OpenStudy (anonymous):

Set 2*sin(x) + 3*cos(x) = C*cos(x+d). Expand the right-hand side and solve for C and d.

myininaya (myininaya):

\[\sin(d)\sin(x)+\cos(d)\cos(x)=\cos(x-d)\] 

myininaya (myininaya):

we want sin(d)=2 and cos(d)=3 so we have